Alyssa Clark is evaluating her debt safety ratio. Her monthly take- home pay is $3,320. Each month, she pays $380 for an auto loan, $120 on a personal line of credit, $60 on a department store charge card, and $85 on her bank credit card. Complete Worksheet 6.1 by listing Alyssa’s outstanding debts, and then calculate her debt safety ratio. Given her current take-home pay, what is the maximum amount of monthly debt payments that Alyssa can have if she wants her debt safety ratio to be 12.5 percent? Given her current monthly debt payment load, what would Alyssa’s take-home pay have to be if she wanted a 12.5 percent debt safety ratio?
1
Expert's answer
2020-05-17T17:55:53-0400

"DSR=\frac{380+120+60+85}{3320}*100=19.43%"

If DRS=12.5% monthly debt payments =X

"12.5=\frac{x}{3320}*100"

"x=\frac{12.5}{100}*3320=415"

If DRS=12.5% take-home pay =y

"12.5=\frac{645}{y}*100"

"y=\frac{645*100}{12.5}=5160"